OpenStudy (anonymous):

HELP ME SOLVE THIS!! (show steps) 3x+5=x√3 HELP ME SOLVE THIS!! (show steps) 3x+5=x√3 @Mathematics

OpenStudy (anonymous):

\[3x + 5 = x \sqrt{3} \] square both sides so u will get, \[9x^2 + 30x + 25 = 3x^2\] \[6x^2 + 30x + 25 = 0\] now i feel u can solve it

OpenStudy (anonymous):

thnx a bunch :)

OpenStudy (anonymous):

its ok buddy

OpenStudy (anonymous):

:D